(1/8+3/10)2+ 1/2*7/8
Answer by CubeyThePenguin(3113)

2(1/8 + 3/10) + (1/2)(7/8)
= 2(17/40) + 7/16
= 17/20 + 7/16
= (17 * 4 + 7 * 5)/80
= 103/80 or 1.2875